OpenStudy (anonymous):

which of the following is a solution for the system y=-3x+5 and y=2x+5? a. (1,7) b.(5,0) c.(2,-1) d.(0,5)

OpenStudy (aroub):

y=-3x+5 y=2x+5 By substitution: 2x+5=-3x+5 2x+3x=-5+5 5x=0 x=0 y=2(0)+5 y=5 => (0,5)

OpenStudy (anonymous):

thankyou!

OpenStudy (aroub):

You're welcome!